OpenStudy (hayleymeyer):

pweezzz haalp ! <333 The volume of a cylindrical can is 15.7 in.3. The diameter is 2 in. Find the height. Use 644-12-03-00-00_files/i0110000.jpg = 3.14. a. 1 in. c. 5 in. b. 62.8 in. d. 3.14 in.

OpenStudy (johnweldon1993):

Hint* \(\large Volume = \pi r^2 h\) So since we are given Volume to be 15.7 And we know the radius is half the diameter so 1/2(2) = 1 We are to use 3.14 for \(\large \pi\) so \[\large 15.7 = 3.14\times 1^2\times h\] and solve for 'h'

OpenStudy (hayleymeyer):

thanks!